A total of $10000 is invested in two funds paying 5% and 7% annual interest. The combined annual interest is $630. How much of the 10000 is invested in each fund?
Answer by checkley75(3666)

.07x+.05(10000-x)=630
.07x+500-.05x=630
.02x=630-500
.02x=130
x=130/.02
x=6500 invested @ 7%.
10000-6500=3500 invested @ 5%
proof
.07*6500+.05*3500=630
455+175=630
630=630